OpenStudy (anonymous):

show all step,,4(x - 6) = -2(x - 3)

OpenStudy (agreene):

4 (x-6) = -2 (x-3) Expand out terms on both sides: 4 x-24 = 6-2 x Add (2 x+24) to both sides: 6 x = 30 Divide both sides by 6: x = 5